# Anomalous diffusion in convergence to effective ergodicity

### v2

@@ Abstract
Power-law exponents in the convergence to effective ergodicity is quantified for Ising-Lenz model in one dimension. Modified Thirumalai-Mountain (TM) metric for magnetisation is computed for the range of temperature values under strongly correlated dynamics. In producing evolution of TM metric over time, time-averaged dynamics is generated by using Metropolis and Glauber single-spin-flip dynamics, and ensemble-averaged dynamics with an exact solution. Superdiffusive behaviour is numerically identified in the parameter regimes studied, i.e., power-law exponents, $\alpha>1.0$.
@@
ergodicity, power-laws, ising models, lattice dynamics, monte carlo
@@ 1 Introduction
Dynamics of cooperation among assembly of independent units has been studied in many different context [1], such as model of magnetic material [2, 3, 4] or the state of a neuron [5]. Convergence to ergodicity in this kind of cooperation dynamics is explored recently [6] using Thirumalai-Mountain (TM) fluctuation metric [7, 8] that is adapted for Ising model’s total magnetization. In this study, we investigate the power-laws emerged from the time evolution of the rate of effective ergodic convergence under the parameters that give rise to strong correlations and anamolous diffusion in covergence. Understanding ergodicity for this circumstances is not only interesting due to fundamental importance in statistical mechanics [9] and anomalous diffusion in general [10], also for its implications in real world applications, such as understanding disruption in neural networks for dementia [11], realisation of associative memory in a solid-state device [12], nature of economic utility [13] and optical lattice dynamics [14].
The Ising model is introduced in Sec. 2. In Sec. 3, the concept of effective ergodicity will be shortly explained. Measuring convergence to ergodicity is shown in Sec. 4. Computation of power-law will be shown in Sec. 5. Summary is given in Sec. 6.
@@ 2 Ising-Lenz System
A lattice with $N$ sites, labelled as $\{s_{i}\}_{i=1}^{N}$, which can take up two values, such as $\{1,-1\}$. These values implies spin up or down in the Ising Model [2, 3, 4] or an activation in neuronal systems [5].
The total energy, Hamiltonian of the system can be written with two interactions, one due to nearest-neighbors (NN) and one due to an external field, with coefficients $J$ and $H$ respectively,
$\mathcal{H}(\{s_{i}\}_{i=1}^{N},J,H)$ $=$ $J\Big((\sum_{i=1}^{N-1}s_{i}s_{i+1})+(s_{1}s_{N})\Big)$ (1)
$+H\sum_{1}^{N}s_{i}.$
the term $s_{1}s_{N}$ is imposed by the periodic boundary conditions which provides translational invariance, making the model a closed chain of interacting units. Thermal scale is expressed with $\beta=\frac{1}{k_{B}T}$ and the corresponding coefficients for NN and external field is scaled by this,
@@
Analytic expression for finite size magnetisation $M_{E}(N,\beta,H)$ is obtained by using Transfer Matrix method. Whereas, time evolution of $M_{E}(N,\beta,H)$ is computed via Monte Carlo procedure with Metropolis and Glauber single-spin-flip dynamics. Where, more details can be found in the previous work [6].
@@ 3 Ergodic Dynamics
A form of ergodic dynamics suggested by Boltzmann that trajectories of many-body system will reach to phase-space partitions where its likely to be in a thermodynamic equilibrium [9]. At this point, ensemble averages and time averages of the system produce close to identical measures. This implies, for a given observable $g$ over a fixed phase-space points, ensemble averaged value can be recovered by time average values, $g(t)$ from $t_{0}$ to $t_{N}$,
@@
where $\langle\rangle$ indicates ensemble averaged value. This kind of basic definition is not standard in the literature [15, 7, 16]. Other forms of ergodicity demands that system should visit all available phase-space partitions, which might not be possible for a finite physical system, and moreover feasibility of this type of ergodicity is questioned recently [17]. In practical terms, since partitions of the phase-space is clustered around few regions, effective ergodicity can be attained quickly [7].
@@
4 Measuring Convergence
@@
Effective ergodic convergence, $\Omega_{G}(t)$ for a given observable $g$, can be constructed based on the fact that identical parts of cooperating units have identical characteristics at thermodynamic equilibrium [7]. This is realised by Thirumalai-Mountain (TM) $G$-fluctuating metric [7, 8], at a given time $t_{k}$
@@
where time average per unit expressed as $g_{j}(t_{k})$ and $\langle g(t_{k})\rangle$ is the instantaneous ensemble average over all units,
@@
where time average per unit expressed as $g_{j}(t_{k})$ and $\langle g(t_{k})\rangle$ is the instantaneous ensemble average over all units,
$g_{j}(t_{k})$ $=$ $\frac{1}{k}\sum_{i=0}^{k}g_{j}(t_{i}),$ (8)
$\langle g(t_{k})\rangle$ $=$ $\frac{1}{N}\sum_{j=1}^{N}g_{j}(t_{k}).$ (9)
Consequently, the rate of ergodic convergence is measured as
where diffusion coeffient expressed as $D_{G}$. Similar measure of ergodicity is used in simple liquids [7, 19], and earthquake fault networks [20, 21].
Using $\Omega_{G}$ for total magnetization in Ising Model at time $t_{k}$ as a function of temperature and external field values reads
@@
where $M_{T}(N,\beta,h)$ and $M_{E}(N,\beta,h)$ correspond to time and ensemble averaged total magnetization. Note that exact expression for $M_{E}$ is used [6]. Hence in computing $\Omega_{M}(t_{k},N,\beta,h)$, we generate Metropolis and Glauber single-spin-flip dynamics for $N=512$ cooperating units, a range of Boltzmann factor values $\beta=[0.5,1.5]$ with a spacing of $0.025$ and non-zero external field values $H=1.0$ with setting short-range interaction strength to $J=1.0$ [22, 23]. Time evolution of inverse rate, $\Omega_{G}^{-1}$ is given in Figure 1a and Figure 1b for $N=512$. After an initial period, inverse rate value decreases rapidly. At higher temperatures, inverse rate behaves differently, we quantify this using power-laws.
@@ 5 Power laws
where $\alpha$ is the the exponent and $t$ is the Monte Carlo time. While, power laws in complex systems and computing them on a given emprical data are studied in depth [24, 25]. We compute power-law exponents of evolution of inverse rate over Monte Carlo time for the parameters given at Figure LABEL:fig:metropolis and LABEL:fig:glauber. Identification where to start to compute $\alpha$ is achived by Kolmogorov-Smirnov test [26]. Temperature dependence of $\alpha$ is reported in Figure 3a and Figure 3b for $N=512$, where $\alpha>1.0$ is observed leading to superdiffusive behaviour. We also observe that there is a sharp increase for $\alpha$ at around $\beta=1.4$, along with a stable value in the range $\beta=(0.8,1.4)$.
6 Summary
Understanding how ergodicity is reached emprically plays a vital role both theoretically and from practical standpoint. We have identified superdiffusive behaviour and power-law exponents for the rate of ergodic convergence in Ising Model under parameters give rise to strong correlation, using the modified Thirumalai-Mountain (TM) metric for the total magnetization. We conclude that combination of stronger temperature values give rise to lower exponents, attributed to high external field.

### v3

@@ Abstract
The nature of diffusion is usually studied for particles or dynamics generating trajectories over time. Similar in principle, these studies can be executed in tracking how a given function of observable properties evolve over time akin to a system’s particle motion, so-called functional-diffusion. This is not the same as systems’ own trajectories but can be considered as a meta-trajectory. Following this idea, we measure how an approach to ergodicity evolves over time for the observed magnetization of a full Ising model with an external field. We compute the functional’s diffusive behavior depending on a range of temperatures via Metropolis and Glauber single-spin-flip dynamics. System’s ensemble-averaged dynamics are computed using expressions from the exact solution. Power-laws on the approach to ergodicity provide the classification of anomalies in the functional-diffusion, demonstrating non-linear super-diffusive behavior over different temperature ranges.
@@
ergodicity, power-laws, Ising models, lattice dynamics, monte carlo, functional-diffusion
@@ 1 Introduction
Brownian motion is probably one of the landmark concepts in statistical physics attracted Einstein’s interest early on [1]. Its importance in formulating statistical mechanics recently reviewed within its 250th anniverssary publications [2]. Major observable in tracing Brownian motion appears as how accumulated displacement curve behaves over time, i.e., linear relationship without intercept corresponding to normal diffusion. Along these lines, dynamics of cooperation among assembly of independent units has been studied in this context [3], such as model of magnetic material [4, 5, 6] and the state of a neuron [7, 8]. Measuring the meta-diffusion in this discrete cases are our core premise in an interesting setting: If the displacement curve of a functional of a function of trajectories shows a power-law scaling over time other than the exponent $1.0$, then the behaviour is called anomalous diffusion [9]. This setting not only yields to mathematically challenging consequences, provides insights into characteristics of the Ising model. Important to note that, described scenario does not apply to trajectories, rather a functional of trajectories. This distinction is the core driver in our work that a functional on the function that explains convergence to ergodicity for total magnetisation is the diffusing, rather than the discrete units of the Ising model.
Convergence to ergodicity in this kind of cooperation dynamics is explored and established [10] using Thirumalai-Mountain (TM) fluctuation metric [11, 12] that is adapted for the Ising model’s total magnetization. In this study, we investigate the power-laws that emerged from the time evolution of the rate of effective ergodic convergence under the parameters that give rise to strong correlations and anomalous diffusion in convergence. Understanding ergodicity for these circumstances is not only interesting due to fundamental importance in statistical mechanics [13] and anomalous diffusion in general but also for its implications in real-world applications, such as understanding disruption in neural networks for dementia [14], realization of associative memory in a solid-state device [15], nature of economic utility [16] and optical lattice dynamics [17].
The formulation of the Ising model is reviewed in Sec. 2. In Sec. 3, the concept of effective ergodicity is explained. Measuring convergence to ergodicity is shown in Sec. 4; this corresponds to the functional-diffusion concept we introduced. Computation of power-law will be shown in Sec. 5. The conclusion is given in Sec. 6.
@@ 2 Ising-Lenz System
A lattice with $N$ sites, labelled as $\{s_{i}\}_{i=1}^{N}$ which can take two values, such as $\{1,-1\}$. These values imply spin up or down in the Ising Model [4, 5, 6] or an activation in neuronal systems [7, 8]. The total energy, the Hamiltonian of the system, can be written with two interactions, one due to nearest-neighbors (NN) and one due to an external field, with coefficients $J$ and $H$ respectively,
$\mathcal{H}({s_{i}}{i=1}^{N},J,H)=J\Big((\sum_{i=1}^{N-1}s_{i}s_{i+1})+(s_{1}s_{N})\Big)+H\sum_{1}^{N}s_{i}.$ (1)
The term $s_{1}s_{N}$ is imposed by the periodic boundary conditions, which provide translational invariance, making the model a closed chain of interacting units. The thermal scale is expressed with $\beta=\frac{1}{k_{B}T}$ and the corresponding coefficients for NN and external field are scaled by this,
@@
The analytic expression for finite size magnetisation $M_{E}(N,\beta,H)$ is obtained by using the Transfer Matrix method. Whereas the time evolution of $M_{E}(N,\beta,H)$ i is computed via the Monte Carlo procedure with Metropolis and Glauber single-spin-flip dynamics. More details can be found in the previous work [10].
@@ 3 Ergodic Dynamics
A form of ergodic dynamics suggested by Boltzmann that trajectories of a many-body system will reach phase-space partitions where it’s likely to be in a thermodynamic equilibrium [13]. At this point, ensemble averages and time averages of the system produce close to identical measures. This implies for a given observable $g$ over a fixed phase-space point, ensemble-averaged value can be recovered by time-averaged values, $g(t)$ from $t_{0}$ to $t_{N}$,
@@
where $\langle\rangle$ indicates ensemble-averaged value. This kind of basic definition is not standard in the literature [18, 11, 19]. Other forms of ergodicity demand that system should visit all available phase-space partitions, which might not be possible for a finite physical system, and moreover feasibility of this type of ergodicity is questioned [20]. In practical terms, since partitions of the phase-space are clustered around few regions, effective ergodicity can be attained quickly [11].
@@
4 Ergodicity convergence as functional-diffusion
@@
Effective ergodic convergence, $\Omega_{G}(t)$ for a given observable $g$, can be constructed based on the fact that identical parts of cooperating units have identical characteristics in thermodynamic equilibrium [11]. This is realised by the Thirumalai-Mountain (TM) $G$-fluctuating metric [11, 12], at a given time $t_{k}$
@@
Functional-diffusion occurs for the development of a function of an observable over time, $F[O(t)]$, measured as its displacement from the initial conditions.In this sense, time-evolution of inverse of $\Omega_{G}(t_{k})$ can be considered as it is in functional-diffusion
where time-averaged per unit expressed as $g_{j}(t_{k})$ and $\langle g(t_{k})\rangle$ is the instantaneous ensemble-averaged over all units,
@@
where diffusion coefficient expressed as $D_{G}$. A similar measure of ergodicity is used in simple liquids [11, 22], and earthquake fault networks [23, 24].
Using $\Omega_{G}$ for the total magnetization in Ising Model at time $t_{k}$ as a function of temperature and external field values is expressed as
@@
where $M_{T}(N,\beta,h)$ and $M_{E}(N,\beta,h)$ correspond to time-averaged and ensemble-averaged total magnetization. Note that exact expression for $M_{E}$ is used [10]. Hence, in computing $\Omega_{M}(t_{k},N,\beta,h)$, we generate Metropolis and Glauber single-spin-flip dynamics for $N=512$ cooperating units, a range of Boltzmann factor values $\beta=[0.5,1.5]$ with a spacing of $0.025$ and non-zero external field values $H=1.0$ with setting short-range interaction strength to $J=1.0$ [25, 26]. The time evolution of inverse rate, $\Omega_{G}^{-1}$, is given in Figure 1a and Figure 1b for $N=512$. After an initial period, inverse rate value decreases rapidly. At higher temperatures, inverse rate behaves differently. Similar results are produced for $N=1024$ on Figure 2a and Figure 2b.
@@ 5 Anomaly in the functional-diffusion
where $\alpha$ is the exponent and $t$ is the Monte Carlo time. While power laws in complex systems and computing them on a given empirical data are studied in depth [27, 28]. We compute power-law exponents of evolution of inverse rate over Monte Carlo time. Identification where to start to compute $\alpha$ is achived by Kolmogorov-Smirnov test [29]. Temperature dependence of $\alpha$ is reported in Figure 3a and Figure 3b for $N=512$, where $\alpha>1.0$ is observed leading to superdiffusive behaviour. We also observe that there is a sharp increase for $\alpha$ at around $\beta=1.4$, along with a stable value in the range $\beta=(0.8,1.4)$. Due to different relaxation regimes, results deviate for $N=1024$, as seen in Figure 4a and Figure 4b, a parabola-like behaviour for superdussion strengths is observed over a temperature dependence.
6 Conclusion
A new concept of functional-diffusion is introduced via a canonical example on Ising model i.e., approach to ergodicity for total magnetisation. Superdiffusive regimes are identified based on different temperature ranges, via computation of power-laws of the inverse fluctuating metric for ergodicity. The results demonstrate size dependence on the relaxation dynamics and their anomalous diffusion for the functional cases. This concept would be a pedagogical test-bed for the extentions of diffusive behaviour beyond particle-trajectories.
Supplement
We provide two notebooks for reproducing the results, data_analysis.ipynb and data_generate.ipynb. Dataset is available on Zenodo [30]